OpenStudy (anonymous):

differentiate y = 8e^x + 2/3^squarerootx

OpenStudy (anonymous):

first part is easy enough since \[\frac{d}{dx}e^x=e^x\]

OpenStudy (anonymous):

so the first part is zero right?

OpenStudy (anonymous):

second part use \[\frac{d}{dx}[b^x]=b^x\ln(b)\]so by the chain rule \[\frac{d}{dx}b^{f(x)}=\ln(b)\times b^{f(x)}\times f'(x)\]

OpenStudy (anonymous):

oh heck no

OpenStudy (anonymous):

the derivative of \(8e^x\) is not \(0\) since \(8e^x\) is not a constant

OpenStudy (anonymous):

the derivative of \(e^x\) is \(e^x\) leave the \(8\) in front

OpenStudy (anonymous):

what do i do about